# Establishing validated standards for Home and Work location Detection

### v1

@@ 2.4.1 Estimating regional employment rates
Figure 4.a shows that both configurations produce similar geographical patterns in regional employment. However, while both exhibit moderate to high Pearson correlation, the maximum configuration has a stronger correlation with reported employment rates but estimates fewer regions due to stricter data coverage thresholds. It also has a higher overall relative error (Figure 4.b, and Figure SI-12, except in Germany, where the trend reverses. This suggests that while the maximum configuration enhances individual accuracy, the minimum configuration may be more suitable for population estimates.
@@ 2.4.2 Home-Work Commuting distances
This case study examines whether commuting distance, measured as the great-circle distance between detected home and work locations, varies with population density and whether parameter configurations can influence results. We estimate urbanization levels and population density around users’ home locations using the 2022 GHS settlement model grid at 1 km resolution [40].
@@
Figure 4: Applications for home and work detection a. Employment rate by province, as reported (left), measured using the configuration with the minimum user loss (middle), and measured with the configuration with the highest accuracy (right). b. Difference between the minimum (light teal) and maximum configurations (dark teal) by country, as the correlation between the reported and selected configuration (top), the relative error (bottom). c. Difference in commuting distance (km) between urban (dashed) and rural (dotted) by configuration. d. Difference in commuting distance (km) against the median and standard error of the population density bins ($\frac{people}{km^{2}}$).
@@ 3 Discussion
In this work, we contributed to the development of accurate and robust methods for home and work detection. Using two datasets (D1, D2) covering $\sim 5,100$ individuals from over $80$ countries, we characterized home/work visit patterns. We found that individuals are most likely at home between midnight and 6 AM ($\sim 75\%$ probability) and at work between 9 AM and 4 PM on workdays ($\sim 50\%$). However, visit patterns and missing data vary widely across users and over time, and even within individual users’ routines.
Building on these insights, we provided an open-source implementation of HoWDe, a home/work detection algorithm that integrates existing approaches while improving robustness against missing data. We validated HoWDe against ground truth datasets, achieving up to $\sim 97\%$ accuracy in D1 and $\sim 88\%$ in D2 for home detection, and up to $\sim 88\%$ accuracy in D1 and $\sim 82\%$ accuracy in D2 for work detection.
Our analysis highlights the trade-off between accuracy and retaining users with low-quality data. We identified two key parameters affecting performance: the window length considered and the minimum visit proportion required for classification. A window of at least $28$ days ensures a reliable home estimation, while work windows longer than $42$ days reduce user retention without significant accuracy gains.
We also revealed how the algorithm’s accuracy varies across population groups, with the largest gaps appearing between urban and suburban areas (for home detection), and between high- and middle-income countries (for work detection). Although these accuracy gaps persist across configurations, two case studies show that prioritizing accuracy over user retention is preferable for individual-level tasks, whereas maximizing retention may be more valuable for population-level analyses.
@@
Despite its strengths, HoWDe has key limitations. While it allows customization of detection windows and thresholds, more advanced methods may be required to fully capture mobility patterns that are collected by diverse GPS-equipped devices [43]. Its performance may also vary with lower temporal resolution or geographically limited datasets, particularly in underrepresented low-income regions with sparser mobile data. Future work should assess its applicability in rural and less densely populated areas [44].
Detecting home and work locations raises important ethical concerns, particularly regarding privacy [45, 46], as the geographical coordinates of these locations are highly identifying and can enable re-identification of individuals in mobility datasets. Currently, many data holders apply in-house heuristics to detect home and work, but these methods are often opaque, introducing methodological uncertainty. By providing a validated and open-source detection framework, our approach enables data holders to perform this sensitive step internally using a standardized and transparent pipeline. We hope this will facilitate the sharing of privacy-preserving datasets—for example, through the release of obfuscated or aggregated representations—while ensuring methodological consistency and scientific utility for downstream research.
@@
4 Material and Methods
4.1 Dataset 1: DTU stop data and self-reported annotations
The first dataset re-uses data collected during 2017-19 already used in previous publications [3], where stop locations were detected using the scalable stop-location detection algorithm InfoStop [47]. The stop location data captures the mobility patterns of 96,065 users between 25 and 64 years old for 81 countries across continents.
The dataset also contains self-recorded home and work labels for $2,309$ labelled home and $2,479$ work locations, collected across $4,599$ mobile phone and smart band users. Note that the labelled dataset is biased towards urban areas and higher- to middle-income countries. The spatial coordinates of locations are no longer available to preserve individuals’ privacy.
All data analysis was carried out in accordance with the European Union’s General Data Protection Regulation 2016/679 (GDPR).
4.2 Dataset 2: Veraset Movements data and annotation
Anonymous raw GPS points are provided at a global level by the Veraset data company in the framework of the World Bank “Monitoring COVID-19 policy response through human mobility data” project. Veraset Movements data consists of raw GPS point information from devices all around the world. In contrast to the majority of GPS data providers, Verasets Movements are sourced from thousands of Software Development Kits (SDK) to reduce sample biases. It features coverage of about $5\%$ of the global population [48].
In our work, unique-device GPS points are processed and analyzed as a proxy of individuals’ mobility and patterns of visits. The data at our disposal spans over a period of $12$ months, starting from Jan 01, 2020, until Dec 31, 2020. In this work, multiple countries were analyzed to assess the impact of the pandemic and the policy responses across a vast spectrum of potentially different behaviors.
Dataset D2, in particular, consists of the GPS trajectories of $100$ distinct devices for each of the $5$ selected countries: Brazil, Colombia, Indonesia, Mexico, and the Philippines. For each country independently, we require $33\%$ of the users to be living in high-wealth neighbourhoods, $33\%$ of the users to be living in medium-wealth neighbourhoods, and the remaining part in low-wealth neighbourhoods. Wealth classification and users’ place of residence are drawn leveraging previous work on GPS mobility in middle-income countries [15]. In particular, each user is also required to have a minimum level of recorded activity, consisting of at least a stop record in $20\%$ in the days before the COVID-19 pandemic declaration and $20\%$ of days with at least a stop record over the entire data period.
This second dataset involves, for each user’s stop location sequences, the expert annotation through detailed inspection of location visits, satellite imagery, and local amenities, focusing on five middle-income countries (Brazil, Colombia, Indonesia, Mexico, and the Philippines) for 500 individuals, from which 287 have at least one work location annotated (see Sec. SI-1 detailed description of the annotation procedure). Given that annotators only labelled the complete longitudinal stop sequences, an explicit consideration of changes in home and work for each individual is not possible.
4.3 Home and Work locations detection algorithms: a primer
Smartphones use an in-built GPS sensor to estimate the device’s latitudes and longitude coordinates over time, at a frequency that varies depending on device usage. The device position is then collected by various location-based apps, such as navigation or social media apps. These data may then be used by location intelligence companies or researchers to study human mobility and its regularities, both at the individual and at the population level. Raw location data utilized for mobility analysis consists of timestamped sequences of (latitude, longitude) records associated with a unique user ID.
The first pre-processing step involved in the analysis is the so-called “stop-location detection” [47, 49]. Stop location detection consists of (i) identifying records that are captured when individuals stay within a small area for a sufficient amount of time (e.g. when they are at home, at the office, or in a restaurant); (ii) applying spatial clustering to all these “static records”, to identify a set of unique locations, i.e. those places where people spend time. After detecting the individuals’ stops-locations, individual trajectories can be summarized as sequences of stays. Each stay is associated with a start time (when the stay started), a duration (how long the stay lasted), and a location ID (where the stay happened).
A second critical processing step often involved in mobility studies is to identify key locations, such as individuals’ home and work locations, starting from the sequences of stays. To achieve this, researchers and practitioners often define heuristics based on the assumption that visits to home and work locations follow specific daily and weekly patterns. While a standard practice does not exist, in this work, we focus on two distinct and widely used state-of-the-art methods that exploit travel circadian rhythms and pattern regularities. These methods were designed and tested over the past decade with different approaches, in a collective scientific effort, separately leveraging diverse sets of mobile-phone-based data (e.g. Call Detail Records (CDRs), eXtended Detail Records, (XDRs), or GPS records) [21, 50, 51, 35] as well as activity-based travel surveys [36]. The main rationale is to rank locations based on the day of the week (weekdays vs. weekends) and the time of day (daytime vs. nighttime) when visits occur.
4.3.1 The Atlas approach
The method used in [2, 6, 52, 28] (hereinafter “Atlas” method) defines a home location as the most probable location visited on weekdays between 10 p.m. and 6 a.m., i.e. the stop location where a device was found spending most of the weekday nighttime. While the Atlas methods mostly focuses on home locations, its definition can be easily generalized to the case of work locations. Following a similar rationale, we define an Atlas work location as the most probable stop location visited during weekdays daytime.
4.3.2 The Time-Geo approach
In contrast with the definition introduced by the Atlas method, more complex heuristics are adopted in Jiang et al. [29] (hereinafter referred to as the “Time-geo” method), in particular for the case of identifying work locations. Time-geo ranks stop locations based on the frequency of visits during weekday nights (between 7 p.m. and 8 a.m.) and weekends and picks the most visited location as an individual’s home. The same logic is applied when detecting the work location, but now focusing on weekday work time boundaries (between 8 a.m. and 7 p.m.) and including additional requirements: i) the candidate’s work location is visited at least three times across the study period, and ii) the distance between the work candidate and an individual’s home location is greater than $500$ meters.
Note that both methods rely on the definition of stop location. Stop location definition might vary depending on the type of data analyzed and the specific characteristics of the data points, e.g. whether you use Call Detail Records (CDR), eXtended Detail Records (XDR), or GPS points. However, the principled choices these algorithms are founded on can naturally be extended to any of these different cases, making them easily generalizable.
In this work, stop locations are defined as spatiotemporally close sets of GPS points. All comparisons between home-work detection methods are always drawn from a set of stop locations constructed following identical procedures. For more information about stop location processing, we refer to the Supplementary Information (SI) Sec.SI-2 and the existing literature [47, 14].
4.4 Home and work visits behavioral clusters
Section 2.1 focused on the temporal attributes of users’ daily mobility patterns. To better highlight groups of users with similar behavioral features, we reported the results in Figure  1.c-f, sorting daily patterns based on the visit hour of the first home visit. behavioral groups were generated using the K-Modes clustering algorithm, a categorical adaptation of the traditional K-Means algorithm [53]. Here, we use daily sequences of 24 hourly bins with three possible stop categories. For example, in the case of home patterns, each sequence consists of either a home (category 1), a visit to “a different location”, e.g., work or recreational locations (category 2), or an hour with missing data (category 3). A similar approach is used for work patterns, where visits to work location are considered as “category 1”, while visits to home location are considered as “a different location” (category 2). The K-Modes algorithm requires as a parameter, the number of desired clusters to compute. In Figure 1.c,e we reported 6 clusters. In Figure 1.d,f we reported 3 and 4 clusters, respectively. Note that, while the number of clusters is the result of the selection of cluster number following the elbow method for each of the four plots, our purpose is to simply highlight the main different visit patterns. Following this idea, in Sec.SI-3, where we present results for the smaller D2 dataset, we manually selected three as the number of clusters used to report the hourly day-visit patterns.
4.5 Behavioral profile entropy
We quantify the degree to which users transition between behavioral profiles measuring the normalized Shannon entropy [38], treating each behavioral cluster as a distinct state, following:
$H_{\text{norm}}=-\frac{1}{\log{k}}\sum_{i=1}^{k}p_{i}\log{p_{i}},$ (1)
where $k$ is the total number of possible clusters, and $p_{i}$ is the probability of occurrence of each state, calculated as: $p_{i}=\frac{n_{i}}{N}$ where $n_{i}$ are the total days classified in cluster $i$ and $N$ the total number of days for a given user. Normalizing by the maximum entropy $\frac{1}{\log{k}}$, ensures that entropy remains comparable across cases (weekend vs. weekday and home vs. work), staying within the same range from 0 (when there is only one profile) to 1 (when there are multiple profiles, and all are equally important).
4.6 HoWDe parameters details
$\mathbf{dhn}$: First, to account for a minimum level of information within a day, we focus only on days with hour sequences consisting of at least three stops, i.e. stops visited in at least three hours of the day (i.e. days with at least three hours of data). The HoWDe algorithm, however, can be tuned to handle different thresholds based on the specific needs or constraints in place.
$\mathbf{dn_{H,W}}$: Second, to account for the potentially varying data quality across different datasets, we introduce two independently configurable pairs of parameters for home and work locations. The first pair of parameters works as a maximum fraction of days without data per sliding window, marking a day with an “undetected home/work location” if the fraction of days without data within the sliding window exceeds the threshold.
$\mathbf{hf_{H,W}}$: A third pair of parameters threshold the minimum fraction of hours per night or day period for a location to be home or work respectively. If the average fraction of visits across all days in the sliding window is below the threshold, the location cannot be considered a potential home/work.
$\mathbf{df_{H,W}}$: Fourth, while home locations are expected to be visited almost every day (provided that enough data are available), work locations are likely to follow a different pattern, as those locations will be subject to local work habits, e.g. four days working week. To account for regularities across days, we introduce an additional requirement for work locations: a work location should be visited for a minimum fraction of days with data.
Table 2: Summary description of the HoWDe configurable parameters. The table reports the parameter names, whether they are used in home detection, work detection, or both, and the description of the parameter purpose. When looking at the hourly bins with data (hours with data), that is the case of $dhn$, $hf_{H}$, and $hf_{W}$, counts and fractions are computed with respect to night/business hourly bins only. Note that, $dhn$, $dn_{H}$, and $dn_{W}$ act as minimum activity thresholds for a given day to have an estimated home or work location estimated. $hf_{H}$, $hf_{W}$, and $df_{W}$ serve a similar purpose but for location-specific visit requirements allowing each location above the threshold to be a candidate home/work location. Specifically, fractions of visits to home/work location candidates are computed relative to the total hours in which an individual has recorded data during the specified period.
Name Usage Description
$\mathbf{\Delta_{T}}$ Home/Work data quality Size of the memory window
$\mathbf{dhn}$ Home/Work data quality Min. number of night/business hourly-bins with data in a day
$\mathbf{dn_{H}}$ Home data quality Max. fraction of days without data in a window
$\mathbf{dn_{W}}$ Work data quality Max. fraction of days without data in a window
$\mathbf{hf_{H}}$ Home estimation Min. fraction of night hourly-bins (avg. over days in the window)
$\mathbf{hf_{W}}$ Work estimation Min. fraction of business hourly-bins (avg. over days in the window)
$\mathbf{df_{W}}$ Work estimation Min. fraction of days within the window
For details on parameter implementation, we refer to the Sec. SI-6, where we provide the algorithm’s pseudocode.
4.7 Accuracy and “fraction of not detected” definition
In this work, we report the performance of the algorithm in terms of two main quantities: “accuracy” and “fraction of not detected”. Here, we provide the definitions of these two concepts. The definitions below are explained in the case of the home detection, and they apply in the same way for the work detection.
The fraction of ‘not detected’ indicates cases where a home location was labelled, but the algorithm couldn’t assign a home due to the thresholds on data quality or minimum fraction of visits. In the case of D1, the dataset consists of a set of user-weeks, corresponding to users that self-reported living in a specific home (the “ground truth home”) in a given week. For some of these user-weeks, our algorithm identifies a “detected home”. We evaluate the performance of the algorithm by comparing pairs of “ground truth home” and “detected home” across user-weeks. The “fraction of not detected” is thus computed as $f_{ND}=1-n_{\bar{H}}/n_{H}$, where $n_{\bar{H}}$ is the number of user-week pairs with a detected home, and $n_{H}$ is the number of user-week pairs with a ground truth home. In the case of D2, where ground truth information is not referring to a specific week, $n_{\bar{H}}$ corresponds to the number of users with a detected home, and $n_{H}$ corresponds to the number of users with a ground truth home.
The “accuracy” captures the fraction of correctly labelled home locations among the set of detected home locations. Therefore, we define accuracy as $Acc=m_{\bar{H}}/n_{\bar{H}}$, where $m_{\bar{H}}$ is the number of user-weeks (for D1) or users (for D2) for which the detected home location matches the ground truth home location. In dataset D2, if annotators identified multiple homes for an individual, we considered it a match if the individual’s detected home location corresponded to at least one of the annotated ground truth home locations. Note that our definition does not correspond exactly to the traditional definition of accuracy because we separately assess the algorithm performance in terms of the detection/non-detection of a home (captured by $f_{ND}$) and the detection of the correct home (captured by $Acc$).
The definitions above are also used in the case of work detection, where only users with a reported work location, $n_{W}$, are included in the denominator of $f_{ND}$, and only those with a detected and a matching work location are counted in the nominator and in the accuracy computation, as $n_{\bar{W}}$ and $m_{\bar{W}}$, respectively.
5 Data and Code Availability
@@
https://doi.org/10.11583/DTU.28846325 [Note: The link will become available upon publication]. To protect user anonymity, we applied the following privacy-preserving transformations to the stop sequences. First, time was discretized into 10-minute intervals. Second, each user’s timeline was shifted so that their first recorded activity occurred in the first week of 1970. Third, entire days were randomly shuffled within each week, while preserving the distinction between weekdays and weekends. All code needed to implement HoWDe is available at [37]. We also provide HoWDe as a Python (PyPI) package at [54].
Acknowledgments
L.L. has been supported by the ERC project “IMMUNE” (Grant agreement ID: 101003183). L.L. acknowledges the support from the “Fondazione Romeo ed Enrica Invernizzi” for the research activities of the ‘Covid Crisis Lab’ at Bocconi University. L.A. and S.D.S. acknowledge support from the Independent Research Fund Denmark (DFF) through the DFF-Research Project 1 (Inge Lehmann) grant “Gender Gaps in Human Mobility”. The authors thank Y. Liao for providing thoughtful feedback on the code release, which helped improve its clarity and usability.
Declaration of Interest statement
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this article.
Author Contributions
S.D.S., L.L., S.F., and L.A. conceived the study. S.D.S., L.L., and O.D.L.C. curated and preprocessed the mobility and ground truth data. Two alternative algorithms were independently developed by L.L., S.F., and O.D.L.C., and by S.D.S. and L.A. The two designs were compared and discussed collaboratively, and the final HoWDe workflow is based on the design by S.D.S. and L.A., with L.L. contributing to its finalization. S.D.S. and L.L. designed the analysis pipeline and conducted the statistical analyses. S.F. and L.A. provided domain expertise and guidance on the interpretation of results. S.D.S. and L.L. drafted the manuscript. All authors discussed the results and approved the final version of the manuscript.

### v2

@@ 6.2.1 Estimating regional employment rates
Figure 5.a shows that both configurations produce similar geographical patterns in regional employment. However, while both exhibit moderate to high Pearson correlation, the maximum configuration has a stronger correlation with reported employment rates but estimates fewer regions due to stricter data coverage thresholds. It also has a higher overall relative error (Figure 5.b, and Supplementary material Figure S-18, except in Germany, where the trend reverses. This suggests that while the maximum configuration enhances individual accuracy, the minimum configuration may be more suitable for population estimates.
@@ 6.2.2 Home-Work Commuting distances
This case study examines whether commuting distance, measured as the great-circle distance between detected home and work locations, varies with population density and whether parameter configurations can influence results. We estimate urbanisation levels and population density around users’ home locations using the 2022 GHS settlement model grid at 1 km resolution [62].
@@
Figure 5: Applications for home and work detection a. Employment rate by province, as reported (left), measured using the configuration with the minimum user loss (middle), and measured with the configuration with the highest detected accuracy (right). b. Difference between the minimum (light teal) and maximum configurations (dark teal) by country, as the correlation between the reported and selected configuration (top), the relative error (bottom). c. Difference in commuting distance (km) between urban (dashed) and rural (dotted) by configuration. d. Difference in commuting distance (km) against the median and standard error of the population density bins ($\frac{people}{km^{2}}$).
@@ 7 Discussion
Home and work location detection from smartphone data is a widely used task in mobility and transport research. Despite its central role in many studies, most existing approaches have not been systematically validated against ground truth and do not explicitly account for the heterogeneous and irregular patterns of missing data that characterise real-world smartphone datasets.
In this work, we proposed HoWDe, an algorithm for home and work location detection from stop-location data. HoWDe integrates key features of state-of-the-art heuristic approaches and improves robustness against missing data by allowing for heterogeneous data gaps within and across users. HoWDe builds on insights derived from empirical home-work visit data. Using two datasets capturing stop-locations and ground truth home and work locations (D1, D2) for $5,099$ individuals from $68$ countries, we characterised home/work visit patterns. We found that individuals are most likely at home between midnight and 6 a.m. ($\sim 75\%$ probability) and at work between 9 a.m. and 4 p.m. on workdays ($\sim 50\%$). However, visit patterns and missing data vary widely across users and over time, and even within individual users’ routines.
We validated HoWDe against the ground truth datasets, achieving up to $\sim 97\%$ detected accuracy in D1 and $\sim 88\%$ in D2 for home detection, and up to $\sim 88\%$ detected accuracy in D1 and $\sim 82\%$ detected accuracy in D2 for work detection. We found that Atlas and Timegeo achieve lower overall detected accuracy than HoWDe. However, they attain high home-detection performance ($81.44\%-82.92\%$ for D1 and $70.10\%-70.62\%$ for D2), indicating that both are adequate for this task. In contrast, for work detection, relying solely on distance and visit counts—as in Timegeo—yields much lower detection accuracy ($37.51\%$ for D1 and $56.89\%$ for D2), whereas incorporating duration, as in our adaptation of Atlas and in HoWDe, results in markedly higher accuracy.
Our analysis highlights the trade-off between achieving accuracy and retaining users with low-quality data. We identified two key parameters affecting performance: the window length considered ($\Delta_{T}$) and the minimum visit proportion required for classification ($f_{hours,H,W}$, $f_{days,W}$). A window of at least $28$ days ensures a reliable home estimation, while work windows longer than $42$ days reduce user retention without significant accuracy gains.
We also revealed how the algorithm’s detected accuracy varies across population groups, with the largest gaps appearing between urban and suburban areas (for home detection), and between high- and middle-income countries (for work detection). Although these accuracy gaps persist across configurations, two case studies show that prioritizing accuracy over user retention is preferable for individual-level tasks, whereas maximizing retention may be more valuable for population-level analyses.
@@
Despite its strengths, HoWDe has key limitations. While providing a robust framework for detecting anchor locations, it does not infer their semantic purpose—that is, it cannot distinguish between work, study, or other activities with similar temporal patterns. Future work could address this by integrating contextual data, such as point-of-interest (POI) information or survey-based semantic labels. In addition, HoWDe assumes a single dominant lifestyle pattern across users and therefore cannot, within a single run, distinguish between regular schedules and rotating night-shift routines. While this assumption aligns with the empirical regularities in our ground truth data, incorporating an initial lifestyle-detection step represents a promising direction for future work. Finally, the current implementation identifies a single home and work location per day and does not explore multi-home or multi-work scenarios. Leveraging the ranked lists of potential home and work locations (stored internally in HoWDe) could support such analyses, but this lies beyond the evaluation possibilities afforded by our ground truth data. More advanced methods may also be needed to handle heterogeneous GPS data and ensure its applicability across contexts, in particular for rural and less densely populated areas [41, 65].
Working with individual-level location data presents well-known challenges for data sharing and reproducibility due to the high risk of re-identification [66, 67]. Under the EU General Data Protection Regulation, re-identification risk arises when: (i) the data is not k-anonymous—that is, individual trajectories are unique—and (ii) the data can be linked, through means that are reasonably likely to be used, to an external dataset containing direct personal identifiers (GDPR Recital 26). The first condition is almost always met in human mobility datasets: even after coarse-graining or obfuscating data points, trajectories remain highly idiosyncratic [66], making k-anonymity infeasible without rendering the data analytically useless. The second condition can be mitigated through privacy-preserving techniques—such as perturbing spatial coordinates, obfuscating timestamps, or omitting sensitive fields—to increase the difficulty of linking mobility traces to known individuals. These transformations can in some cases reduce re-identification risk to a negligible level, but often at the cost of limiting the usefulness of the data for certain analytical purposes. This creates a fundamental trade-off between privacy protection and scientific utility, which is why access to individual-level location data is typically restricted to secure environments and regulated by data-use agreements.
Because of these constraints, many data holders perform sensitive inference steps—such as detecting home and work locations—internally. However, the methods used for this purpose are often proprietary or undocumented, introducing uncertainties and reducing reproducibility. Our contribution directly addresses this gap by providing a validated, open-source pipeline for detecting home and work locations that can be executed in-house by data holders prior to data sharing. This supports methodological transparency and consistency, while enabling the release of privacy-preserving data products—such as aggregated visitation patterns or synthetic traces.
@@
8 Code and Data Availability
All code needed to implement HoWDe is available as a GitHub repository at [61], along with usage examples in the GitHub folder /tutorials. We additionally provide HoWDe as a Python package on PyPI for direct installation via pip.
@@ 8 Code and Data Availability
https://doi.org/10.11583/DTU.28846325. To protect user anonymity, we applied the following privacy-preserving transformations to the stop sequences. First, time was discretised into 10-minute intervals. Second, each user’s timeline was shifted so that their first recorded activity occurred in the first week of 1970. Third, entire days were randomly shuffled within each week, while preserving the distinction between weekdays and weekends. All data analysis was carried out in accordance with the European Union’s General Data Protection Regulation 2016/679 (GDPR).